ADNAN MUHAMMED

Grade 12,

If x= 2-√5/2+√5 and y= 2+√5/2-√5, find the value of (x*x)-(y*y)

If x= 2-√5/2+√5 and y= 2+√5/2-√5, find the value of (x*x)-(y*y)

Grade:9

1 Answers

Arun
6 years ago
Dear student
 
x = 2-√5/2+√5 
on rationalization
 
x = (4 + 5 – 4 √5) / – 1
 
x =   4 √5 – 9
 
Now similarly y =   – 4 √5 – 9
 
Now question is asked for x^2 – y^2
 = (x+y) * (x – y)
 
 =  – 18 * 8 √5
 = – 144√5